Problem 2 Consider the following source code, where R, S, and T are constants declared with #define. long int AIRISJIT int store_ele(int h, int i, int j, long int dest) Alh][i]01-.dest; return sizeof(A) In compiling this program, GCC generates the following assembly code (with-02) store_ ele: movslq %esi, %rsi movslq %edi, %rdi movq (%rcx), %rax leaq (%rdi,%rdi,4), %rdi leaq (%rsi,%rsi,4), %rcx movslq %edx, %rdx addq %rcx, %rdx movq %rax, A( %rdx,8) movl $1120, %eax ret Create a text file to show your work and the steps required to determine, R, S, and T Use your reverse engineering skills to determine the values of R, S, and T, based upon the assembly code Be sure to show all the calculations required to determine R, S, and T. (15 points each for R. S. T, 15 points for steps/calculations)
Answer #1

For an array declared as Type DIRI[C], array element DO) is at address where L is the size of the data type in bytes. A. It fmovl 12(%ebp), %edx #moves j to %edx leal (%edx, %edx, 4), %eax #multiplies %edx times 5, stores in %eax leal (%edx, %eax, 2)xd = A L=4 (R(9% + j) + k) R=11 s=9 (99x + 11j + k) = (11(9x + j) + k) Since it returns 1980 as sizeof, we know the size of t